OpenStudy (anonymous):

rewrite in radian measure. -268 degrees

OpenStudy (anonymous):

multiply with pi/180

OpenStudy (tkhunny):

\(\dfrac{\pi}{180º}\)

OpenStudy (luigi0210):

\[\huge \mathbb{\color{green}{WELCOME~TO~OPENSTUDY}}\]